# Mixture of Experts Made Personalized: Federated Prompt Learning for Vision- Language Models

### v1

@@ Abstract
Prompt learning for pre-trained Vision-Language Models (VLMs) like CLIP has demonstrated potent applicability across diverse downstream tasks. This lightweight approach has quickly gained traction from federated learning (FL) researchers who seek to efficiently adapt VLMs to heterogeneous scenarios. However, current federated prompt learning methods are habitually restricted to the traditional FL paradigm, where the participating clients are generally only allowed to download a single globally aggregated model from the server. While justifiable for training full-sized models under federated settings, in this work, we argue that this paradigm is ill-suited for lightweight prompts. By facilitating the clients to download multiple pre-aggregated prompts as fixed non-local experts, we propose Personalized Federated Mixture of Adaptive Prompts (pFedMoAP), a novel FL framework that personalizes the prompt learning process through the lens of Mixture of Experts (MoE). pFedMoAP implements a local attention-based gating network that learns to generate enhanced text features for better alignment with local image data on the client, benefiting from both local and downloaded non-local adaptive prompt experts. The non-local experts are sparsely selected from a server-maintained pool, fostering collaborative learning across clients.To evaluate the proposed algorithm, we conduct extensive experiments across 9 datasets under various heterogeneous federated settings. The results show that pFedMoAP consistently outperforms the state-of-the-art alternatives, underscoring its efficacy in personalizing prompt learning for CLIP within the federated learning paradigm.
@@ 2 Related work
Personalized federated learning. The mediocre performance of conventional federated learning (FL) (McMahan et al., 2017) over heterogeneous data calls for a more customized solution. Personalized federated learning (PFL) (Tan et al., 2022; Kulkarni et al., 2020), allowing tailored models for each client instead of a consensus model, systemically mitigates the data heterogeneity issue. Various strategies have been proposed to achieve personalization, including local fine-tuning (Cheng et al., 2021), different optimization techniques such as attentive message passing and regularization (Huang et al., 2021; Li et al., 2021a), and different personalized layers including feature extractor, classification head, attention, and batch normalization layers(Liang et al., 2020; Arivazhagan et al., 2019; Sun et al., 2023; Li et al., 2021b). Some recent PFL works also switch gear to model heterogeneity (Xie et al., 2024; Yi et al., 2024a). While promising in addressing data heterogeneity, these methods primarily focus on traditional machine learning models and do not leverage the capabilities of large pre-trained vision-language models.
@@
Federated prompt learning for pre-trained models. Prompt learning for pre-trained models combines models’ generalization capabilities and prompt learning’s flexibility in adapting these models to downstream tasks. As a fundamental work for this combination, CoOp (Zhou et al., 2022b) models a prompt’s context words with learnable vectors in CLIP. Such a combination has quickly drawn the attention of the FL community. For instance, PromptFL (Guo et al., 2023b) does a FedAvg (McMahan et al., 2017) style global aggregation over the locally updated prompts. Building upon this, FedPR (Feng et al., 2023) learns federated visual prompts for MRI reconstruction, while Fed-DPT (Su et al., 2022) applies both visual and textual prompt tuning to facilitate domain adaptation over decentralized data. pFedprompt (Guo et al., 2023a) integrating global consensus prompt and local attention over stored few-shot data for each client. FedOTP (Li et al., 2024) and FedPGP (Cui et al., 2024) employ Optimal Transport and prompt-wise contrastive loss, respectively, between global and local prompts, capturing diverse category traits on a per-client basis. However, these methods habitually adhere to the restriction for locally learned prompts from inter-client sharing before aggregation. For full-sized models, this inter-client sharing is evidently prohibitive, as it introduces $O(N^{2}\cdot M)$ communication cost per round if each of the $N$ models with size $M$ are shared to $N$ clients (Luo & Wu, 2022; Luo et al., 2023). For the prompt, however, $M$ can be orders of magnitude smaller (see Fig. 1), which substantially lowers the communication cost and frees federated prompt learning from this ill-suited restriction.
@@
Federated learning with Mixture of Experts. From the proprietary GPT-4 (Achiam et al., 2023) to the open-sourced Mixtral of Experts 8$\times$7B (Jiang et al., 2024), Qwen1.5-MoE (Bai et al., 2023), Mixture of Experts (MoE) (Zhou et al., 2022c; Masoudnia & Ebrahimpour, 2014) is prevailing long after its initial proposal thanks to the recent heat in Large Language Models (LLMs). Its application in FL, however, dates back before LLMs. FedMix (Reisser et al., 2021) and FedJETs (Dun et al., 2023) allow each client to construct an MoE with a shared gating network, selecting specific non-local models more adaptive to the client’s local data for ensembling. PFL-MoE(Guo et al., 2021) and pFedMoE Yi et al. (2024b) incorporate the global aggregated model as the global expert and its locally fine-tuned model as a local expert to each client, achieving a personalized two-expert MoE. However, under federated settings, with too many experts on the client comes prohibitively high communication overhead (Reisser et al., 2021; Dun et al., 2023) while too few experts impairs the benefit of employing MoE.
@@ 4.3 Attention-based gating network: in pFedMoAPand beyond
As such, our design of the attention-based gating network extends beyond federated learning. This versatile network could be adapted to general prompt learning for VLMs where multiple prompts are trained, such as multi-task learning or domain adaptation. We anticipate that our work will inspire future research exploring the broader applications of mixture of adaptive prompts for VLMs. The detailed pFedMoAP algorithm can be found in Algorithm 1.
@@ 5.1 Experimental setup
Table 1: Few-shot performance on CLIP datasets under pathological non-IID setting. The highest accuracy for each dataset is in bold.
@@
Implementation details. 1) CLIP datasets. Each dataset in CLIP datasets is partitioned into $N=10$ clients, each with a disjoint set of classes evenly and randomly assigned to the clients. The training proceeds for $T=10$ rounds with $r=100\%$ participation rate. The CLIP uses a ResNet50 backbone. 2) CIFAR10 & CIFAR100. $N=10$ clients results from $\text{Dir}(\alpha=0.5)$ partition with $T=120$ rounds with $r=10\%$. CLIP also uses a ResNet50 backbone. 3) DomainNet & Office-Caltech10. Each domain of these two datasets is partitioned to 5 clients with $\text{Dir}(\alpha=0.3)$, resulting in $N=30$ for DomainNet and $N=20$ for Office-Caltech10. $T=25$ for both datasets, while $r=25\%$ for DomainNet and $r=50\%$ for Office-Caltech10. The CLIP uses a ViT-b-16 backbone. 4) Training specs. For all methods, we use SGD as the optimizer with a learning rate of 0.002 and 5 local epochs (except CoOp is locally trained for 25 epochs without FL). For pFedMoAP, we use SGD with a learning rate of 0.01 to train the $h=8$-head gating network and default $\lambda=0.5$ in Eq. 10. The core source code of the pFedMoAP is included in the supplementary material, and the full implementation will be made publicly available upon acceptance of the paper.
Table 2: Results on DomainNet with feature shift and label shift with $\text{Dir}(\alpha=0.3)$ partition into 5 clients/domain. The highest accuracy for each domain is in bold.
Clipart Infograph Painting Quickdraw Real Sketch Average
@@ 5.2 Performance evaluation
Table 3: Results on CIFAR10 and CIFAR100 with label shift with (Dir($\alpha=0.5$)) partition into 100 clients. The highest accuracies are in bold.
@@
Table 4: Results on Office-Caltech10 with feature shift and label shift with $\text{Dir}(\alpha=0.3)$ partition into 5 clients/domain. The highest accuracy for each domain is in bold.
@@
Evaluation on Feature & Label Shifts. To evaluate the performance of our method in scenarios closer to real-world FL applications, we introduce feature shift on top of label shift via DomainNet and Office-Caltech10where each domain is partitioned into 5 clients with $\text{Dir}(\alpha=0.3)$, resulting in 30 and 20 total clients in DomainNet and Office-Caltech10, respectively. The results of these are presented in Tabs. 2 and 4, respectively. Under two types of heterogeneity, traditional federated learning methods struggle to keep benefiting the clients. On average, the FedAvg style method PromptFL exhibits worse performance than the local training method CoOp by 12.51% and 8.96% respectively on DomainNet and Office-Caltech10. However, pFedMoAP remains better performance over both local training and FL with up to 5.94% boost in accuracy under two types of heterogeneity. This validates the effectiveness and robustness of pFedMoAP in scenarios closer to real-world federated settings.
@@ 6 Conclusion
In this paper, we introduced pFedMoAP, a novel framework for personalized federated prompt learning in CLIP-like Vision-Language Models. Our approach leverages the lightweight nature of prompts to enable efficient cross-client knowledge sharing through a Mixture of Experts paradigm. By implementing a client-specific, attention-based gating network, pFedMoAP dynamically incorporates both local and non-local expert knowledge, striking a balance between personalization and collaborative learning. Our extensive experiments across various datasets and federated settings demonstrate the superior performance of pFedMoAP compared to state-of-the-art alternatives, particularly in handling extreme data heterogeneity. The consistent outperformance of pFedMoAP in scenarios with both label and feature shifts underscores its robustness and efficacy in real-world federated learning applications. Furthermore, our ablation studies highlight the method’s adaptability to different shot numbers and its stability across various hyperparameter settings. As federated learning continues to evolve, pFedMoAP represents a significant step forward in personalizing prompt learning for VLMs, opening new avenues for efficient and effective collaborative learning in privacy-preserving settings.

### v3

@@ Abstract
Federated prompt learning benefits federated learning with CLIP-like Vision-Language Model’s (VLM’s) robust representation learning ability through prompt learning. However, current federated prompt learning methods are habitually restricted to the traditional FL paradigm, where the participating clients are generally only allowed to download a single globally aggregated model from the server. While justifiable for training full-sized models under federated settings, in this work, we argue that this paradigm is ill-suited for lightweight prompts. By facilitating the clients to download multiple pre-aggregated prompts as fixed non-local experts, we propose Personalized Federated Mixture of Adaptive Prompts (pFedMoAP), a novel FL framework that personalizes the prompt learning process through the lens of Mixture of Experts (MoE). pFedMoAP implements a local attention-based gating network that learns to generate enhanced text features for better alignment with local image data, benefiting from both local and downloaded non-local adaptive prompt experts. Extensive experiments on 9 datasets under various federated settings demonstrate the efficacy of the proposed pFedMoAP algorithm. The code is available at https://github.com/ljaiverson/pFedMoAP.
@@ 2 Related work
Personalized federated learning. The mediocre performance of conventional federated learning (FL) (McMahan et al., 2017) over heterogeneous data calls for a more customized solution. Personalized federated learning (PFL) (Tan et al., 2022; Kulkarni et al., 2020), allowing tailored models for each client instead of a consensus model, systemically mitigates the data heterogeneity issue. Various strategies have been proposed to achieve personalization, including local fine-tuning (Cheng et al., 2021), different optimization techniques such as attentive message passing and regularization (Huang et al., 2021; Li et al., 2021a), and different personalized layers including feature extractor, classification head, attention, and batch normalization layers (Liang et al., 2020; Arivazhagan et al., 2019; Sun et al., 2023; Li et al., 2023; 2021b). Some works use separate models to learn global and local knowledge, respectively (Cai et al., 2024; Deng et al., 2020). Some recent PFL works also switch gear to model heterogeneity (Xie et al., 2024; Yi et al., 2024a). While promising in addressing data heterogeneity, these methods primarily focus on traditional machine learning models and do not leverage the capabilities of large pre-trained vision-language models.
@@
Federated prompt learning for pre-trained models. Prompt learning for pre-trained models combines models’ generalization capabilities and prompt learning’s flexibility in adapting these models to downstream tasks. As a fundamental work for this combination, CoOp (Zhou et al., 2022b) models a prompt’s context words with learnable vectors in CLIP. Such a combination has quickly drawn the attention of the FL community. For instance, PromptFL (Guo et al., 2023b) does a FedAvg (McMahan et al., 2017) style global aggregation over the locally updated prompts. Building upon this, FedPR (Feng et al., 2023) learns federated visual prompts for MRI reconstruction, while Fed-DPT (Su et al., 2022) applies both visual and textual prompt tuning to facilitate domain adaptation over decentralized data. pFedprompt (Guo et al., 2023a) integrating global consensus prompt and local attention over stored few-shot data for each client. FedOTP (Li et al., 2024) and FedPGP (Cui et al., 2024) employ Optimal Transport and prompt-wise contrastive loss, respectively, between global and local prompts, capturing diverse category traits on a per-client basis. Pan et al. (2024) presents a theoretical analysis framework for prompt-based federated learning of vision-language models. However, these approaches habitually adhere to the paradigm where locally learned prompts are not allowed for inter-client sharing before aggregation. While sharing full models between clients would be impractical due to the massive communication costs – $O(N^{2}\cdot M)$ per round when $N$ models of size $M$ are distributed to $N$ clients (Luo & Wu, 2022; Luo et al., 2023) – sharing prompts is a different scenario. Since prompts are orders of magnitude smaller than full models (as shown in Fig. 1), the communication overhead becomes much more affordable. This makes it feasible for the inter-client sharing of pre-aggregated prompts under federated settings.
@@
Federated learning with Mixture of Experts. From the proprietary GPT-4 (Achiam et al., 2023) to the open-sourced Mixtral of Experts 8$\times$7B (Jiang et al., 2024), Qwen1.5-MoE (Bai et al., 2023), Mixture of Experts (MoE) (Zhou et al., 2022c; Masoudnia & Ebrahimpour, 2014) is prevailing long after its initial proposal thanks to the recent heat in Large Language Models (LLMs). Its application in FL, however, dates back before LLMs. FedMix (Reisser et al., 2021) and FedJETs (Dun et al., 2023) allow each client to construct an MoE with a shared gating network, selecting specific non-local models more adaptive to the client’s local data for ensembling. PFL-MoE(Guo et al., 2021) and pFedMoE (Yi et al., 2024b) incorporate the global aggregated model as the global expert and its locally fine-tuned model as a local expert to each client, achieving a personalized two-expert MoE. However, under federated settings, with too many experts on the client comes prohibitively high communication overhead (Reisser et al., 2021; Dun et al., 2023) while too few experts impairs the benefit of employing MoE.
@@ 4.3 Attention-based gating network: in pFedMoAP and beyond
As such, our design of the attention-based gating network extends beyond federated learning. This versatile network could be adapted to general prompt learning for VLMs where multiple prompts are trained, such as multi-task learning or domain adaptation. We anticipate that our work will inspire future research exploring the broader applications of mixture of adaptive prompts for VLMs. The detailed pFedMoAP algorithm is described at Algorithm 1.
@@ 5.1 Experimental setup
Table 1: Few-shot performance on CLIP datasets under pathological non-IID setting.
@@
Implementation details. 1) CLIP datasets. Each dataset in CLIP datasets is partitioned into $N=10$ clients, each with a disjoint set of classes evenly and randomly assigned to the clients. The training proceeds for $T=10$ rounds with $r=100\%$ participation rate. The CLIP uses a ResNet50 backbone. 2) CIFAR10 & CIFAR100. $N=100$ clients results from $\text{Dir}(\alpha=0.5)$ partition with $T=120$ rounds with $r=10\%$. CLIP also uses a ResNet50 backbone. 3) DomainNet & Office-Caltech10. Each domain of these two datasets is partitioned to 5 clients with $\text{Dir}(\alpha=0.3)$, resulting in $N=30$ for DomainNet and $N=20$ for Office-Caltech10. $T=25$ for both datasets, while $r=25\%$ for DomainNet and $r=50\%$ for Office-Caltech10. The CLIP uses a ViT-b-16 backbone. 4) Training specs. For all methods, we use SGD as the optimizer with a learning rate of 0.002 and 5 local epochs (except CoOp is locally trained for 25 epochs without FL). For pFedMoAP, we use SGD with a learning rate of 0.01 to train the $h=8$-head gating network and default $\lambda=0.5$ in Eq. 10. A summary of the dataset setup can be found at Sec. A.1.
@@ 5.2 Performance evaluation
Table 2: Results on CIFAR10 and CIFAR100 with label shift with (Dir($\alpha=0.5$)) partition into 100 clients
@@
Table 3: Results on DomainNet with feature shift and label shift with $\text{Dir}(\alpha=0.3)$ partition into 5 clients/domain
Clipart Infograph Painting Quickdraw Real Sketch Average
Table 4: Results on Office-Caltech10 with feature shift and label shift with $\text{Dir}(\alpha=0.3)$ partition into 5 clients/domain
@@
Evaluation on Feature & Label Shifts. To evaluate the performance of our method in scenarios closer to real-world FL applications, we introduce feature shift on top of label shift via DomainNet and Office-Caltech10 where each domain is partitioned into 5 clients with $\text{Dir}(\alpha=0.3)$, resulting in 30 and 20 total clients in DomainNet and Office-Caltech10, respectively. The results of these are presented in Tabs. 3 and 4, respectively. Under two types of heterogeneity, traditional federated learning methods struggle to benefit the clients. On average, the FedAvg style method PromptFL exhibits worse performance than the local training method CoOp by 12.51% and 8.96% respectively on DomainNet and Office-Caltech10. However, pFedMoAP remains better than local training and FL with up to 5.94% accuracy boost under both types of heterogeneity. This validates the effectiveness and robustness of pFedMoAP in scenarios closer to real-world federated settings.
5.3 Ablation study
@@ 6 Conclusion
In this paper, we introduced pFedMoAP, a novel framework for personalized federated prompt learning in CLIP-like VLMs. Our approach leverages the lightweight nature of prompts to enable efficient cross-client knowledge sharing through a Mixture of Experts paradigm. By implementing a client-specific, attention-based gating network, pFedMoAP dynamically incorporates both local and non-local expert knowledge, striking a balance between personalization and collaborative learning. Extensive experiments across various datasets and federated settings demonstrate the superior performance of pFedMoAP compared to state-of-the-art alternatives, particularly in handling extreme data heterogeneity with label and/or feature shift. As federated learning continues to evolve, pFedMoAP represents a significant step forward in personalizing prompt learning for VLMs, opening new avenues for efficient and effective collaborative learning in privacy-preserving settings.
@@ Acknowledgments
APPENDIX
Appendix A Additional experiments
A.1 Details of dataset setup
We follow the most recent prompt-based FL for VLMs works’ standard (Li et al., 2024; Cui et al., 2024) and use the 9 datasets they have used. Tab. 5 lists the details of these datasets and the partitioning method used in our experiments for each dataset.
Table 5: The statistics of each dataset and the partitioning details used in our experiments
Dataset Training set size Test set size #Classes #Clients Sample Rate Heterogeneity
Flowers102 4,093 2,463 102 10 100% Pathological
OxfordPets 2,944 3,669 37 10 100% Pathological
Food101 50,500 30,300 101 10 100% Pathological
Caltech101 4,128 2,465 100 10 100% Pathological
DTD 2,820 1,692 47 10 100% Pathological
Office-Caltech10 2,025 508 10 20 50% Dir(0.3)
DomainNet 18,278 4,573 10 30 25% Dir(0.3)
CIFAR10 50,000 10,000 10 100 10% Dir(0.5)
CIFAR100 50,000 10,000 100 100 10% Dir(0.5)
A.2 pFedMoAP under differential privacy
In pFedMoAP, clients remain unaware of the origin of non-local prompts, including which client they come from or the training round they were last updated in. Besides, the non-local prompts are downloaded after a KNN-based sparse selection process on the server. These factors collectively make it extremely difficult for a client to infer the gradient associated with a prompt or the data it has been trained with, which largely mitigates the privacy risk.
Table 6: Performance under ($\epsilon$, $\delta$)-differential privacy on CLIP datasets under pathological non-IID setting.
Flowers102 OxfordPets Food101 Caltech101 DTD
Without differential privacy (from Tab. 1)
PromptFL+FedProx (Li et al., 2020b) 66.40$\pm$0.29 89.24$\pm$0.41 76.24$\pm$1.94 89.41$\pm$0.55 44.26$\pm$1.11
pFedMoAP(ours) 98.41$\pm$0.04 99.06$\pm$0.09 93.39$\pm$0.09 97.95$\pm$0.07 89.13$\pm$0.54
With differential privacy ($\epsilon=50$)
PromptFL (Guo et al., 2023b) 67.07$\pm$0.60 88.05$\pm$0.32 77.41$\pm$0.60 84.83$\pm$0.42 38.39$\pm$1.25
PromptFL+FedProx (Li et al., 2020b) 66.22$\pm$0.63 87.78$\pm$0.61 77.27$\pm$0.59 84.68$\pm$0.64 39.43$\pm$1.11
pFedMoAP(ours) 98.34$\pm$0.06 99.08$\pm$0.02 93.36$\pm$0.04 97.90$\pm$0.08 89.99$\pm$0.49
With differential privacy ($\epsilon=25$)
PromptFL (Guo et al., 2023b) 64.25$\pm$1.10 86.26$\pm$1.07 76.84$\pm$0.66 85.00$\pm$1.59 38.19$\pm$0.66
PromptFL+FedProx (Li et al., 2020b) 62.87$\pm$0.99 86.82$\pm$0.47 76.21$\pm$0.64 84.51$\pm$1.52 37.82$\pm$0.52
pFedMoAP(ours) 98.36$\pm$0.12 99.02$\pm$0.04 93.41$\pm$0.13 97.99$\pm$0.06 89.11$\pm$0.28
Additionally, we evaluate how ($\epsilon$, $\delta$)-differential privacy (DP) (Dwork, 2006) affects the performance of pFedMoAP on CLIP datasets and compare its performance with PromptFL and PromptFL+FedProx. This is achieved by clipping the gradients of the prompts and adding noise to the uploaded prompts. We set the failure probability $\delta$ to be $0.05$ and report the accuracy with the privacy budget $\epsilon$ being set to $50$ and $25$. The results in Tab. 6 demonstrate that while the accuracy of PromptFL and PromptFL+FedProx significantly drop with privacy guarantees, DP’s impact in accuracy performance to pFedMoAP remains minimal for different privacy budgets. This is because the gating network in pFedMoAP is maintained locally for the entirety of the training process, and its multi-head attention layer learns proper transformations for the experts to align with the image features, mitigating the impact of the noise-contaminated prompts.
A.3 Attention-based vs. linear projection-base gating network
Table 7: Comparison between the proposed attention-based gating network and the linear projection-based gating network
Flowers102 OxfordPets Food101 Caltech101 DTD
Linear projection-based (3 experts) 86.92$\pm$1.84 90.54$\pm$1.33 78.19$\pm$3.07 89.59$\pm$1.46 61.42$\pm$5.43
Linear projection-based (10 experts) 69.64$\pm$4.57 52.78$\pm$6.88 77.39$\pm$3.29 86.57$\pm$1.96 30.42$\pm$7.14
Attention-based, with aggregation 97.56$\pm$0.07 98.24$\pm$0.12 91.89$\pm$0.19 96.17$\pm$0.18 87.52$\pm$0.69
Attention-based, without aggregation (ours) 98.41$\pm$0.04 99.06$\pm$0.09 93.39$\pm$0.09 97.95$\pm$0.07 89.13$\pm$0.54
We compare the proposed attention-based gating network against the traditional linear projection-based gating network. The results are reported in Tab. 7 which strongly favor the attention-based approach, showing significantly better and more stable performance across all datasets. Notably, the linear projection-based approach struggles with larger numbers of experts (10 experts vs 3 experts), while the attention-based method maintains high performance. This is because the attention-based gating network 1) is more robust to adaptive experts; 2) serves as linear probing with more capacity; 3) leverages CLIP’s feature alignment with attention mechanism; and 4) is agnostic to experts’ order (detailed strength of attention-based approach over linear projection is described in Sec. 4.3).
In addition, we choose to avoid aggregating the gating network due to its larger size. The gating network can be magnitudes larger than the prompts (see Tab. 9), which would largely increase the communication overhead. This design choice also aims for higher performance. As the deepest parameterized module of the entire model, the gating network should be fully “personalized” as opposed to “globalized” to achieve higher performance (Arivazhagan et al., 2019; Collins et al., 2021). The process of aggregating the gating network limits the level of its personalization, which yields worse results than our method that keeps the gating network locally (see Tab. 7).
A.4 Additional ablation study
Table 8: Ablation study: impact of number of experts ($K$ non-local experts + $1$ local experts) to pFedMoAP on DomainNet
Clipart Infograph Painting Quickdraw Real Sketch Average
$K+1=5$ 47.20$\pm$0.30 46.80$\pm$0.48 32.54$\pm$0.42 37.67$\pm$0.52 31.50$\pm$0.65 36.09$\pm$0.92 38.63$\pm$0.29
$K+1=10$ 46.89$\pm$1.07 46.15$\pm$0.95 32.76$\pm$0.42 37.70$\pm$1.02 31.94$\pm$0.83 36.84$\pm$1.21 38.71$\pm$0.45
$K+1=15$ 47.36$\pm$1.12 46.29$\pm$0.83 32.76$\pm$0.57 37.76$\pm$0.98 31.78$\pm$0.72 36.67$\pm$1.03 38.77$\pm$0.38
$K+1=20$ 47.49$\pm$0.64 46.73$\pm$0.71 32.74$\pm$0.84 37.16$\pm$0.34 31.02$\pm$0.59 37.67$\pm$0.72 38.80$\pm$0.11
$K+1=25$ 47.56$\pm$0.92 46.89$\pm$1.30 32.23$\pm$0.97 38.03$\pm$1.14 31.66$\pm$0.77 36.79$\pm$1.02 38.86$\pm$0.43
$K+1=30$ 47.56$\pm$0.87 46.64$\pm$1.38 32.19$\pm$0.89 37.93$\pm$1.22 31.80$\pm$0.95 37.01$\pm$1.08 38.85$\pm$0.43
Number of experts. We present an ablation study examining how the number of experts ($K$ non-local experts $+$ 1 local expert) affects pFedMoAP’s performance on the DomainNet dataset. The results in Tab. 8 show that the average performance gradually improves as the number of experts increases from 5 to 30. Fig. 5 plots the trend of performance with different numbers of experts from 5 to 40 on CIFAR10. For both datasets, it plateaus around larger numbers of experts, suggesting that while having multiple experts is beneficial, there’s a point of diminishing returns. This is because the extra downloaded experts (selected last by our KNN-based expert selection mechanism) by a client are trained from the data that is distinctly distributed from the client’s local data distribution. This shows that the proposed method does not require an excessive number of experts to achieve high performance.
Figure 5: The impact of the number of experts on CIFAR10 with 100 clients
Table 9: Ablation study: the impact of output dimension of CLIP backbone to pFedMoAP on CLIP datasets. For $d_{\text{feature}}<1024$, a pooling layer is added after the $d_{\text{feature}}=1024$ feature from the backbone to reduce the size of the gating network as mentioned in Sec. 4.3
Gating network size Flowers102 OxfordPets Food101 Caltech101 DTD
$d_{\text{feature}}=32$ 4.2K 97.28$\pm$0.18 98.75$\pm$0.32 93.42$\pm$0.08 97.37$\pm$0.08 88.61$\pm$0.89
$d_{\text{feature}}=64$ 16.6K 98.55$\pm$0.10 98.91$\pm$0.23 93.89$\pm$0.12 97.75$\pm$0.12 89.96$\pm$0.09
$d_{\text{feature}}=128$ 66.0K 98.41$\pm$0.04 99.06$\pm$0.09 93.39$\pm$0.09 97.95$\pm$0.07 89.13$\pm$0.54
$d_{\text{feature}}=256$ 263.2K 99.01$\pm$0.05 98.88$\pm$0.21 92.49$\pm$0.20 97.93$\pm$0.07 90.88$\pm$0.16
$d_{\text{feature}}=512$ 1.1M 98.18$\pm$0.38 96.85$\pm$0.22 90.34$\pm$0.31 96.99$\pm$0.11 89.65$\pm$0.10
$d_{\text{feature}}=1024$ 4.2M 98.11$\pm$0.33 95.81$\pm$0.84 89.20$\pm$0.37 96.82$\pm$0.26 89.03$\pm$0.14
(a) Caltech101, 10 experts
(b) DTD, 10 experts
(c) Food101, 10 experts
(d) Flowers102, 10 experts
(e) OxfordPets, 10 experts
(f) DomainNet, 20 experts
Figure 6: Contribution of the experts based on averaged attention score across all test images. The first five charts are for CLIP datasets, for which there are 10 clients in each dataset. The last chart is for DomainNet with a total of 30 clients.
Feature dimension. We investigate the impact of the output dimension ($d_{\text{feature}}$) of the CLIP backbone on pFedMoAP’s performance across CLIP datasets. The study shows that moderate feature dimensions ($128$-$256$) generally yield the best results, with $d_{\text{feature}}=128$ or $256$ consistently performing well across all datasets. Lower feature dimensions ($d_{\text{feature}}=32$ or $64$) have less capacity, failing to unlock the full potential of the pre-trained VLM. High dimensions ($d_{\text{feature}}=128$ or $1024$) significantly increase the size of the gating network. With less training data on a client, high dimensions lead to overparameterized gating network, resulting in slightly degraded performance. This suggests that a relatively compact feature dimension is sufficient and possibly preferable for the attention-based gating mechanism.
Visualization of contribution for each expert. To have a clear understanding of how the experts contribute to the MoE text feature (${\bm{T}}_{MoE}$), we visualize the contribution of the local and non-local experts towards it. This is achieved by taking the following steps. For each image in a client’s test set, we compute the attention score (softmax of scaled dot product between the image and text features (Vaswani, 2017)) for each expert (local and non-local) through the multi-head attention layer in the proposed gating network. The scores from non-local experts are sorted in a descending manner. All scores are averaged over the test set across all clients. We visualize the averaged scores under 6 datasets in Fig. 6. From this figure, we can see that 1) every expert has a non-trivial contribution to ${\bm{T}}_{MoE}$, the text feature generated from the MoE, indicating that each expert is indispensable in the whole design; 2) as a result of the minimization, the contribution from the local experts to the MoE is not the highest. This is reasonable since besides the ${\bm{T}}_{MoE}$, Eq. 10’s second term optimizes the text feature from the local expert itself (${\bm{T}}_{L}$) based on the image feature, which directly contributes to the loss; and 3) although not the highest, the contribution from the local expert to the MoE is still among the highest ones as it is the only expert that is updated during local training with fixed non-local prompts.